a I of 9.65A is passed through 100 ml NaCl for 100 s. what will be the ph of resulting solution?assume 100% current efficiency and constancy of volume of solution

Grade:12th pass

1 Answers

Vikas TU
3 years ago
Given current = 9.65 A
Time = 1000 s
Charge carried Q = It
= 9.65 A(1000 s)
= 9,650 C
9650 C× \frac{1mol  e^{-}}{96485 C}   × \frac{2 mol NaOH}{2 mol  e^{-}}   × \frac{40 g NaOH}{1 mol NaOH}  
= 4.0 g NaOH